OpenStudy (anonymous):

Check my answer - C Which of the following statements best describes the graph of x + y = 4? It is a line which intersects the x-axis at (4, 4). It is a line which intersects the y-axis at (4, 4). It is a line joining the points whose x- and y-coordinates add up to 8. It is a line joining the points whose x- and y-coordinates add up to 4.

OpenStudy (cderaalien):

Your answer is wrong, I believe.

OpenStudy (cderaalien):

Usually when you have an equation like this, it will have lines that intersect.

OpenStudy (cderaalien):

Have you ever heard of that?

OpenStudy (anonymous):

@CderaAlien um im not sure

OpenStudy (cderaalien):

Here is an example.

OpenStudy (cderaalien):

The point where the lines intersect at is the solution to the equation.

OpenStudy (cderaalien):

Do you understand what I mean?

OpenStudy (cderaalien):

The reason why your answer is wrong is, because you added 4...and you can't do that with this type of equation.

OpenStudy (cderaalien):

At least, not when it's asking how it's graphed.

OpenStudy (anonymous):

Oh! i see..

OpenStudy (skullpatrol):

Which of the following statements best describes the graph of x + y = 4? It is a line joining the points whose x- and y-coordinates add up to 4.

OpenStudy (cderaalien):

You can't give direct answers. ^^

OpenStudy (cderaalien):

You have to try to help them understand @skullpatrol

OpenStudy (skullpatrol):

I agree, @CderaAlien. But the question is tricky enough to justify letting the OP see how the answer is deduced from the question, imho.

OpenStudy (cderaalien):

Okay, I can't say I don't understand. Fair enough.

OpenStudy (skullpatrol):

@queencel draw a table of x values and calculate the corresponding y values from the given equation. $$\huge x + y = 4$$

OpenStudy (skullpatrol):

Let x = 0, 1, 2, ... Then plot the points on a coordinate plane. :-)

OpenStudy (skullpatrol):

That^ will help you understand why the other answer choices are wrong.

OpenStudy (anonymous):

mhm got it!

OpenStudy (anonymous):

Actually C is correct!

OpenStudy (skullpatrol):

C says the x- and y-coordinates add up to 8? But the equation says x + y = 4.

OpenStudy (skullpatrol):